22 Probability (Mathematics) questions asked in BPSC TRE previous year papers, each with the correct answer and explanation. Read them free on UTpaper, or attempt the full papers as timed online tests.
A bag contains 2 red, 3 green and 2 blue balls. Two balls are drawn at random. The probability that none of the balls drawn is blue isएक थैले में 2 लाल, 3 हरी तथा 2 नीली गेंदें हैं। दोनों गेंदों को यादृच्छिक रूप से निकाला जाता है। दोनों गेंदों में से किसी के भी नीला न होने की प्रायिकता है
A10/2110/21
B11/2111/21
C2/72/7
DMore than one of the aboveउपर्युक्त में से एक से अधिक
A bag contains 5 red balls and some blue balls. If the probability of drawing a blue ball from the bag is four times that of a red ball, then the number of blue balls in the bag isएक थैले में 5 लाल एवं कुछ नीली गेंदें हैं। यदि एक नीली गेंद निकालने की प्रायिकता, लाल गेंद निकालने की प्रायिकता की चार गुनी है, तब थैले में स्थित नीली गेंदों की संख्या है
A1010
B4040
C2020
DMore than one of the aboveउपर्युक्त में से एक से अधिक
ENone of the aboveउपर्युक्त में से कोई नहीं
Show answerHide answer
Correct answer: (C)2020
Explanation
P(blue) = 4 × P(red) means the blue balls are 4 times the 5 red balls = 20.P(नीली) = 4 × P(लाल), अतः नीली गेंदें लाल (5) की 4 गुनी = 20 हैं।
A carton consists of 100 shirts of which 88 are good, 8 have minor defects and 4 have major defects. Jimmy, a trader, will only accept the shirts which are good, but Sujata, another trader, will only reject the shirts which have major defects. One shirt is drawn at random from the carton. The proba- bilities that it is not acceptable to Jimmy and acceptable to Sujata areएक बॉक्स में 100 शर्ट हैं, जिनमें 88 अच्छी हैं, 8 में कम दोष है तथा 4 में बड़ा दोष है। जिम्मी एक व्यवसायी है जो केवल अच्छी शर्ट स्वीकार करता है परन्तु सुजाता एक अन्य व्यवसायी है जो केवल उन शर्टों को अस्वीकार करती है जिनमें बड़ा दोष है। एक शर्ट को यादृच्छया बॉक्स से निकाला जाता है। जिम्मी द्वारा इसे स्वीकार न करने एवं सुजाता द्वारा इसे स्वीकार करने की प्रायिकताएँ हैं
A0.12, 0.960.12, 0.96
B0.88, 0.960.88, 0.96
C0.12, 0.040.12, 0.04
DMore than one of the aboveउपर्युक्त में से एक से अधिक
ENone of the aboveउपर्युक्त में से कोई नहीं
Show answerHide answer
Correct answer: (A)0.12, 0.960.12, 0.96
Explanation
P(not acceptable to Jimmy) = (8 + 4)/100 = 0.12; P(acceptable to Sujata) = 96/100 = 0.96.P(जिम्मी को अस्वीकार्य) = (8 + 4)/100 = 0.12; P(सुजाता को स्वीकार्य) = 96/100 = 0.96।
Two different dice are rolled together. Find the probability of getting the sum of numbers on two dice to be 5.दो अलग-अलग पासों को एकसाथ उछाला जाता है। दोनों पासों पर आने वाली संख्याओं का योग 5 होने की प्रायिकता ज्ञात कीजिए।
A1/91/9
B1/41/4
C1/61/6
DMore than one of the aboveउपर्युक्त में से एक से अधिक
ENone of the aboveउपर्युक्त में से कोई नहीं
Show answerHide answer
Correct answer: (A)1/91/9
Explanation
Sum 5 occurs for (1,4), (2,3), (3,2), (4,1): 4/36 = 1/9.योग 5 के लिए (1,4), (2,3), (3,2), (4,1): 4/36 = 1/9।
One ticket is drawn at random from a bag containing tickets numbered 1 to 40. The probability that the selected ticket has a number, which is a multiple of 7, is1 से 40 तक की संख्या वाले टिकटों से भरे एक बैग से यादृच्छिक रूप से एक टिकट निकाला जाता है। चयनित टिकट की संख्या 7 का गुणक होने की प्रायिकता है
A1/71/7
B1/81/8
C7/407/40
DMore than one of the aboveउपर्युक्त में से एक से अधिक
ENone of the aboveउपर्युक्त में से कोई नहीं
Show answerHide answer
Correct answer: (B)1/81/8
Explanation
Multiples of 7 up to 40 are 7, 14, 21, 28, 35 (5 numbers), so P = 5/40 = 1/8.40 तक 7 के गुणज 7, 14, 21, 28, 35 (5 संख्याएँ) हैं, अतः P = 5/40 = 1/8।
In a lottery, there are 6 prizes and 24 blanks. What is the probability of not getting a prize?एक लॉटरी में, 6 पुरस्कार और 24 रिक्त-स्थान हैं। पुरस्कार न मिलने की प्रायिकता क्या है?
A3/43/4
B3/53/5
C4/54/5
DMore than one of the aboveउपर्युक्त में से एक से अधिक
Q8BPSC TRE 2.0 (2023) Mathematics and Science Class 6-8
One card is drawn from a well-shuffled deck of 52 cards. If each outcome is equally likely, then what is the probability that the card will not be a diamond?अच्छी तरह से फेंटी गई 52 ताशों की एक गड्डी में से एक ताश खींचा जाता है। यदि प्रत्येक परिणाम सम सम्भाव्य हो, तो इसकी प्रायिकता क्या है कि ताश ईंट (diamond) नहीं होगा?
A1/41/4
B12/1312/13
C3/43/4
DMore than one of the aboveउपर्युक्त में से एक से अधिक
ENone of the aboveउपर्युक्त में से कोई नहीं
Show answerHide answer
Correct answer: (C)3/43/4
Explanation
There are 13 diamonds, so P(not a diamond) = 39/52 = 3/4.ईंट के 13 पत्ते हैं, इसलिए P(ईंट नहीं) = 39/52 = 3/4।
Q9BPSC TRE 2.0 (2023) Mathematics and Science Class 6-8
Two cards are randomly drawn from a pack of 52 cards and thrown away. What is the probability of drawing an ace in a simple draw from the remaining 50 cards?52 ताशों की एक गड्डी में से दो ताश यादृच्छया निकाले जाते हैं और फेंक दिए जाते हैं। शेष 50 ताशों में से एक सरल निकासी में एक इक्का निकलने की प्रायिकता क्या है?
A7/507/50
B11/5011/50
C9/509/50
DMore than one of the aboveउपर्युक्त में से एक से अधिक
ENone of the aboveउपर्युक्त में से कोई नहीं
Show answerHide answer
Correct answer: (C)9/509/50
Explanation
Note: the official key gives (C) 9/50; however, since the two discarded cards are random, the probability of an ace in the next draw is still 4/52 = 1/13, which is not among the options. Official key kept.ध्यान दें: आधिकारिक कुंजी (C) 9/50 है; परन्तु फेंके गए दो ताश यादृच्छिक होने से इक्का मिलने की प्रायिकता अब भी 4/52 = 1/13 रहती है, जो विकल्पों में नहीं है। आधिकारिक कुंजी रखी गई है।
Q10BPSC TRE 2.0 (2023) Mathematics Class 9-10 (Re-exam)
A box contains 10 mangoes out of which 4 are rotten. 2 mangoes are taken out together. If one of them is found to be good, then probability that the other is also good isएक डिब्बे में 10 आम हैं जिनमें से 4 सड़े हुए हैं। 2 आम एक साथ निकाले जाते हैं। यदि उनमें से एक अच्छा पाया जाता है, तो संभावना यह कि दूसरा भी अच्छा है, क्या है?
A1/31/3
B8/158/15
C5/135/13
DMore than one of the aboveउपर्युक्त में से एक से अधिक
ENone of the aboveउपर्युक्त में से कोई नहीं
Show answerHide answer
Correct answer: (C)5/135/13
Explanation
Given at least one good: P(both good) = C(6,2)/C(10,2) = 15/45 and P(at least one good) = 1 − C(4,2)/45 = 39/45. The conditional probability is 15/39 = 5/13.कम से कम एक अच्छा होने पर: P(दोनों अच्छे) = C(6,2)/C(10,2) = 15/45 और P(कम से कम एक अच्छा) = 1 − C(4,2)/45 = 39/45। सप्रतिबंध प्रायिकता 15/39 = 5/13 है।
Q11BPSC TRE 2.0 (2023) Mathematics Class 9-10 (Re-exam)
For a given probability distribution
f(x) = (1/8) C(3, x), x = 0, 1, 2, 3
for a random variable X, the moment generating function of this isकिसी दिए गए संभाव्यता वितरण
f(x) = (1/8) C(3, x), x = 0, 1, 2, 3
के लिए यादृच्छिक चर X का आघूर्ण-जनक फलन है
A(1 + eᵗ)²(1 + eᵗ)²
B(1/4) eᵗ(1/4) eᵗ
C(1/8)(1 + eᵗ)³(1/8)(1 + eᵗ)³
DMore than one of the aboveउपर्युक्त में से एक से अधिक
Q12BPSC TRE 2.0 (2023) Mathematics Class 9-10 (Re-exam)
What is the second moment about the origin for a random variable X with probability density?
f(x) = x/2, 0 < x < 2; 0 otherwiseनिम्न संभाव्यता घनत्व वाले यादृच्छिक चर X के मूल के बारे में दूसरा आघूर्ण क्या है?
f(x) = x/2, 0 < x < 2; अन्यथा 0
A22
B33
C44
DMore than one of the aboveउपर्युक्त में से एक से अधिक
If a coin is tossed twice, the probability of getting at best one head isएक सिक्के को दो बार फेंककर कम से कम एक सिर प्राप्त करने की प्रायिकता है
A1/41/4
B3/43/4
C1/21/2
DMore than one of the aboveउपर्युक्त में से एक से अधिक
ENone of the aboveउपर्युक्त में से कोई नहीं
Show answerHide answer
Correct answer: (B)3/43/4
Explanation
Outcomes HH, HT, TH, TT. At most one head: HT, TH, TT = 3/4 (at least one head is also 3/4).परिणाम HH, HT, TH, TT; अधिकतम/कम से कम एक सिर: 3 परिणाम, प्रायिकता 3/4।
A husband and wife appear in an interview for two vacancies in the same post. The probability of husband’s selection is 1/7 and that of wife’s selection is 1/5. Then the probability that both of them are selected isदो रिक्तियों के एक पोस्ट के साक्षात्कार में एक पति और पत्नी उपस्थित रहे। अगर पति के चयन की संभावना 1/7 और पत्नी की 1/5 है, तो वे दोनों की चयन की संभावना है
A12/3512/35
B1/351/35
C11/3511/35
DMore than one of the aboveउपर्युक्त में से एक से अधिक
The probability that A solves a problem is 1/2 and that of B is 2/3. The probability that the problem is solved isA द्वारा किसी प्रोब्लेम को हल करने की प्रायिकता 1/2 और B की 2/3 है, तो प्रोब्लेम हल हो जाने की प्रायिकता है
A1/31/3
B1/61/6
C5/65/6
DMore than one of the aboveउपर्युक्त में से एक से अधिक
What is the probability that a leap year, selected at random, will have 53 Sundays?यादृच्छया चुने गए किसी अधिवर्ष में 53 रविवार होने की प्रायिकता क्या है?
A3/73/7
B5/75/7
C2/72/7
DMore than one of the aboveउपर्युक्त में से एक से अधिक
ENone of the aboveउपर्युक्त में से कोई नहीं
Show answerHide answer
Correct answer: (C)2/72/7
Explanation
A leap year has 52 weeks and 2 extra days. Of the 7 equally likely pairs of consecutive days, 2 contain a Sunday (Sat–Sun, Sun–Mon), so the probability is 2/7.अधिवर्ष में 52 सप्ताह और 2 अतिरिक्त दिन होते हैं। लगातार दिनों के 7 समसम्भाव्य जोड़ों में से 2 में रविवार आता है (शनि–रवि, रवि–सोम), इसलिए प्रायिकता 2/7 है।
There are two children in a family. If it is known that at least one of the children is a boy, what is the probability that both the children are boys?एक परिवार में दो बच्चे हैं। यदि यह ज्ञात हो कि बच्चों में से कम-से-कम एक बच्चा लड़का है, तो दोनों बच्चों के लड़का होने की क्या प्रायिकता है?
A1/41/4
B1/31/3
C3/43/4
DMore than one of the aboveउपर्युक्त में से एक से अधिक
ENone of the aboveउपर्युक्त में से कोई नहीं
Show answerHide answer
Correct answer: (B)1/31/3
Explanation
The equally likely cases with at least one boy are BB, BG and GB. Only BB has two boys, so the probability is 1/3.कम-से-कम एक लड़के वाली समसम्भाव्य स्थितियाँ BB, BG और GB हैं। इनमें केवल BB में दोनों लड़के हैं, इसलिए प्रायिकता 1/3 है।
From a group of 7 men and 4 women, a committee of 6 persons is formed. The probability that the committee consists of exactly 2 women is7 पुरुषों और 4 महिलाओं के एक समूह से 6 व्यक्तियों की एक समिति बनाई गई। समिति में ठीक 2 महिलाएँ होने की प्रायिकता है
A5/115/11
B4/114/11
C7/117/11
DMore than one of the aboveउपर्युक्त में से एक से अधिक
A pair of dice is thrown then the probability of getting 7 as the sum, isदो पासों को युग्मतः फेंका जाता है, तब उनका योग 7 आने की प्रायिकता है
A1/61/6
B5/65/6
C11
DMore than one of the aboveउपर्युक्त में से एक से अधिक
ENone of the aboveउपर्युक्त में से कोई नहीं
Show answerHide answer
Correct answer: (A)1/61/6
Explanation
There are 36 equally likely outcomes, and a sum of 7 comes from (1,6), (2,5), (3,4), (4,3), (5,2), (6,1), i.e. 6 outcomes. Probability = 6/36 = 1/6.36 समसम्भाव्य परिणाम हैं, और योग 7 (1,6), (2,5), (3,4), (4,3), (5,2), (6,1) से आता है, अर्थात् 6 परिणाम। प्रायिकता = 6/36 = 1/6।
The probability that a leap year will have 53 Fridays, isएक अधिवर्ष में 53 शुक्रवार होने की प्रायिकता है
A1/71/7
B2/72/7
C3/73/7
DMore than one of the aboveउपर्युक्त में से एक से अधिक
ENone of the aboveउपर्युक्त में से कोई नहीं
Show answerHide answer
Correct answer: (B)2/72/7
Explanation
A leap year has 366 days = 52 weeks + 2 days. The 2 extra days can be any of 7 consecutive pairs (Sun-Mon, ..., Sat-Sun), and Friday occurs in 2 of them (Thu-Fri, Fri-Sat). So P = 2/7.अधिवर्ष में 366 दिन = 52 सप्ताह + 2 दिन होते हैं। 2 अतिरिक्त दिन लगातार दिनों के 7 जोड़ों (रवि-सोम, ..., शनि-रवि) में से कोई भी हो सकते हैं, और शुक्रवार 2 जोड़ों (गुरु-शुक्र, शुक्र-शनि) में आता है। अतः P = 2/7।
How to use these BPSC TRE Mathematics questions
Every question here was asked in a real BPSC TRE paper; the paper and shift are shown on the card, and the number opens the question on its own page. Try each one before tapping Show answer, then attempt the full BPSC TRE papers as timed tests to check your speed under exam conditions.